Suppose you work at an organic milk company, your boss asked you to evaluate a project with an infinite life. Sales and costs project to $1,000 and $500 per year, respectively. (Assume sales and costs occur at the end of the year [i.e., profit of $500 at the end of year one]). There is no depreciation and the tax rate is 21 percent. The real required rate of return is 10 percent. The inflation rate is 4 percent and is expected to be 4 percent forever. Sales and costs will increase at the rate of inflation. If the project costs $3,000, what is the NPV?

a) $950.00

b) $1,629.62

c) $365.38

d) $472.22
Answer #1

Inflation, sales and cost increase at same percentage so given values are all real values of sales and cost per year

Net profit = Sales - cost = 1000 - 500 = 500

Tax = 21% * 500 = 0.21 * 500 = 105

After tax cash flow = 500 - 105 = 395

Present value of this infinite sereis = 395 / 0.10 = 3950

NPP of the project = -3000 + 3950 = 950

First option is correct